OpenStudy (anonymous):

f(x)=3x^2+1 find f(x+3)-[f(x)+3] help me solve this please.

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

i know that we start of with f(x+3)=3(x+3)^2+1

OpenStudy (anonymous):

but everytime i expand its not right

OpenStudy (anonymous):

f(x+3)=3(x+3)^2 now put the value of f(x) and f(x+3) as asked in ques u will get the ans

OpenStudy (anonymous):

yes the last one i have the correct answer to

OpenStudy (anonymous):

and the first one it's like i can only write it out but when i expand i mess it up

OpenStudy (anonymous):

i should get 18x +24 for final answer

hero (hero):

Find f(x + 3): f(x+3) = 3(x+3)^2 + 1 = 3(x^2 + 6x + 9) + 1 = 3x^2 + 18x + 27 + 1 = 3x^2 + 18x + 28 Then find f(x) + 3: f(x) + 3 = 3x^2 + 1 + 3 = 3x^2 + 4 so f(x + 3) - (f(x) + 3) = (3x^2 + 18x + 28) - (3x^2 + 4) Finish it..

OpenStudy (anonymous):

(3x^2-3x^2) +12x(28-4) 18x+24

hero (hero):

f(x+ 3) - (f(x) + 3) = (3x^2 - 3x^2) + 18x + (28 - 4) = 18x - 24

hero (hero):

12x(28 - 4) is not a correct expression.

OpenStudy (anonymous):

(3x^2-3x^2) +18x+(28-4)

hero (hero):

Yes, that is correct

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

final will be 18x-24

OpenStudy (anonymous):

thanks for helping me @hero